# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 1.3. Explicit symmetry breaking and the emergence of a unique Kähler–Einstein metric
Next, we break the $\mathcal{G}$–symmetry explicitly by fixing a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and a $\mathcal{K}$–invariant metric on $K_{(X,\Delta)}$ with positive curvature. This metric defines a $\mathcal{K}$–invariant Kähler form $\omega_{0}$ in $c_{1}(-K_{(X,\Delta)})$ and thus a symplectic form on $X.$ The natural action of $\mathcal{K}$ on $K_{(X,\Delta)}$ gives rise to a moment map
@@
Given a positive integer $N$ such that 0 is a regular value of $\boldsymbol{m}_{N}$ we denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 6.1 for the general definition. The algebraic definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 6.11), i.e. that
(1.13) $\mu_{\boldsymbol{m}_{N}}^{(N)}\coloneqq\frac{1}{Z_{\boldsymbol{m}_{N}}}\nu_{\boldsymbol{m}_{N}}^{(N)},\,\,\,\,Z_{\boldsymbol{m}_{N}}\coloneqq\int\nu_{\boldsymbol{m}_{N}}^{(N)}$
defines a probability measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}.$ The probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ is symmetric, i.e. invariant under permutations of $X^{N}$ and $\mathcal{K}$–invariant. The original $\mathcal{G}$–symmetry has thus been reduced to a $\mathcal{K}$–symmetry. Accordingly, the probability measure 1.13 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e. $\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.13.
Before stating our two main conjectures we note that in contrast to the $\mathcal{K}$–reduced probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)},$ the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ depends on a choice of bases in $H^{0}(X,-kK_{(X,\Delta)}).$ As a consequence, so does its total mass $Z_{\boldsymbol{m}_{N}}.$ In what follows we will take the basis to be orthonormal with respect to the $L^{2}$–norm on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the fixed metric on $K_{(X,\Delta)}.$ The corresponding constant $Z_{\boldsymbol{m}_{N}}$ shall then be called the $\mathcal{K}$-reduced partition function, in view of its statistical mechanical interpretation described in Section 6.1. The fixed metric on $K_{(X,\Delta)}$ also determines the Mabuchi functional $\mathcal{M}$ on the space of Kähler metrics in $c_{1}(-K_{(X,\Delta)})$ (which, a priori, is only defined up to an additive constant).
@@
(1.14) $\text{(i)}-\lim_{N\rightarrow\infty}N^{-1}\log Z_{\boldsymbol{m}_{N}}=\inf\mathcal{M}.$
As a consequence, $Z_{\boldsymbol{m}_{N}}<\infty$ for $N$ sufficiently large and viewing the empirical measure $\delta_{N}$ as a random variable on $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right),$
@@
(1.16) $(\delta_{N})_{*}\nu_{\boldsymbol{m}_{N}}^{(N)}\sim e^{-NF(\mu)},$
where $F$ is the free energy functional, introduced in [10], that may be identified with the Mabuchi functional on the space of all Kähler metrics $\omega$ in $c_{1}(-K_{(X,\Delta)})$ via the relation $F(\omega^{n}/V)=\mathcal{M}(\omega)$. We show that the validity of such a LDP would imply both the asymptotics 1.14 and the convergence 1.15, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$. Moreover, by introducing temperature into the picture the corresponding LDP would also imply Conjecture 1.2, using the observation that the infimum over $\mathcal{G}$ in formula 1.8 is attained at the unique Kähler metric in a given $\mathcal{G}$–orbit whose volume form has vanishing moment (see Prop 3.4).
Our next main result establishes the first part of the previous conjecture when $n=1$ with an explicit rate of convergence:
@@
The convergence 1.14 holds on any K-polystable log Fano curve $(X,\Delta)$ (i.e. when $n=1$), when $N$ is taken as the sequence of odd positive integers. More precisely,
$N^{-1}\log Z_{\boldsymbol{m}_{N}}=\inf\mathcal{M}+O\left(\frac{\log N}{N}\right).$
As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.15) we will show, in the companion paper [4], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 6.2).
@@
where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.17 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [60]. Such integrals can be computed explicitly in terms of Gamma values, as first shown in [53] in the context of Conformal Field Theory, and letting $N\rightarrow\infty$ allows us to deduce Theorem 1.7, by exploiting some connections to K-stable arithmetic log Fano curves in [3].
It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.17 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 8, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [3], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [3], is a logarithmic generalization of Odaka’s modular height [86]. The proof of Theorem 1.7 reveals that this arithmetic conjecture holds for all K-polystable log Fano curves $(X,\Delta)$ defined over the integers.
@@ 1.5. Outlook on valuative aspects
where $X^{\text{div}}$ denotes the space of all divisorial valuations $v$ over $X,$ $A_{(X,\Delta)}(v)$ denotes the log discrepancy of $v$ (formula 2.6), $\left\|\mu\right\|$ denotes the non-Archimedean pluricomplex energy of a probability measure $\mu$ on $X^{\text{div}}$ with respect to $K_{(X,\Delta)}$ [35] and $\mathcal{P}(X^{\text{div}})_{\text{ss}}$ denotes the space of all probability measures $\mu$ on $X^{\text{div}}$ of the following form (where $N$ is the sequence defined by formula 1.6):
@@
The convergence in the second item above is motivated by the approach for proving the convergence of $\gamma^{(N)}(X,\Delta)$ towards $\delta(X,\Delta)_{\text{ }}$ outlined in [17, Section 5], using non-Archimedean pluripotential theory. The validity of the third item would imply the first item, by [48]. It should be emphasized that, in general, the invariant $\delta(X,\Delta)_{\text{ps }}$ differs from the reduced $\delta$-invariant of $(X,\Delta),$ introduced in [93], which is defined with respect to a fixed maximal complex torus in $\mathcal{G}.$ Our results imply Conjecture 1.8 for log Fano curves, but we leave the proof and further investigations of Conjecture 1.8 for the future.
@@ 1.6. Acknowledgements
This work was supported by a Wallenberg Scholar grant from the Knut and Alice Wallenberg foundation.
@@ 1.7. Organization of the paper
  • In Section 4 applications of Theorem 3.13 to constrained logarithmic Hardy–Littlewood–Sobolev inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
  • In Section 5 we establish a sharp version of the constrained logarithmic Hardy–Littlewood–Sobolev inequalities on the two-sphere, by combining Theorem 3.13 with the uniqueness result for the Liouville equation on $S^{2}$ established in [72].
  • In Section 6 we study the $\mathcal{K}$–reduced probability measures $\mu_{\boldsymbol{m}_{N}}^{(N)}$ on $X^{N}$ and the corresponding partition function $Z_{\boldsymbol{m}_{N}}$ appearing in Conjecture 1.6. The asymptotics as $N\rightarrow\infty$ are described in a overarching conjectural Large Deviation Principle, formulated in Conjecture 6.4, that we show implies both Conjecture 1.1 and Conjecture 1.2.
  • In Section 7 Theorem 1.7 is established, showing that $N^{-1}\log Z_{\boldsymbol{m}_{N}}$ converges towards the inf of the Mabuchi functional on any log Fano curve with vanishing Futaki invariant.
@@ 2.1. Setup
We start by describing in more detail the general setup, introduced in Section 2. Let $(X,\Delta)$ be a log Fano manifold, i.e. $X$ is a complex projective manifold and $\Delta$ is an effective $\mathbb{Q}$–divisor on $X$ such that the log anticanonical line bundle $K_{(X,\Delta)}$ of $(X,\Delta)$ is ample. Recall that $K_{(X,\Delta)}$ is the $\mathbb{Q}$–line bundle defined as the tensor power of the canonical line bundle $K_{X}$ of $X$ and the $\mathbb{Q}$–line bundle defined by $\Delta,$ i.e. $K_{(X,\Delta)}\coloneqq K_{X}+\Delta,$ using the standard additive notation for tensor powers of line bundles that we shall adopt. Fix a positive integer $k$ such that $k\Delta$ is a divisor – rather than a mere $\mathbb{Q}$–divisor – so that $kK_{(X,\Delta)}$ defines a line bundle and set
@@
$\text{Supp}\left(\mathcal{D}^{(N)}\right)\coloneqq\left\{(x_{1},\dots,x_{N})\colon\exists s\in H^{0}(X,-kK_{(X,\Delta)})-\{0\}\text{ such that }s(x_{1})=\cdots=s(x_{N})\right\}$

### v3

@@ 1.5. Explicit symmetry breaking and the emergence of a unique Kähler–Einstein metric
Let $(X,\Delta)$ be a log Fano manifold and fix a maximally compact subgroup $\mathcal{K}$ of the corresponding symmetry group $\mathcal{G}$ and a $\mathcal{K}$–invariant metric on $K_{(X,\Delta)}$ with positive curvature. The curvature defines a Kähler form $\omega_{0}$ in $c_{1}(-K_{(X,\Delta)})$ and thus a $\mathcal{K}$–invariant symplectic form on $X.$ As consequence, the natural action of $\mathcal{K}$ on $K_{(X,\Delta)}$ gives rise to a moment map
@@
Consider first the case when 0 is a regular value of $\boldsymbol{m}_{N}.$ We then denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 6.2 for the general definition. The algebro-geometric definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 6.13), i.e. that
(1.15) $\mu_{0}^{(N)}\coloneqq\frac{1}{Z_{N,0}}\nu_{0}^{(N)},\,\,\,\,Z_{N,0}\coloneqq\int\nu_{0}^{(N)}$
defines a probability measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}.$ The probability measure $\mu_{0}^{(N)}$ is symmetric, i.e. invariant under permutations of $X^{N}$ and $\mathcal{K}$–invariant, but not $\mathcal{G}$-invariant. This is an instance of explicit symmetry breaking, as the original $\mathcal{G}$–symmetry has been reduced to a $\mathcal{K}$–symmetry. Accordingly, the probability measure 1.15 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{0}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e. $\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.15.
Before stating our two main conjectures we note that in contrast to the $\mathcal{K}$–reduced probability measure $\mu_{0}^{(N)},$ the measure $\nu_{0}^{(N)}$ depends on a choice of bases in $H^{0}(X,-kK_{(X,\Delta)}).$ As a consequence, so does its total mass $Z_{N,0}.$ In what follows we will take the basis to be orthonormal with respect to the $L^{2}$–norm on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the fixed metric on $K_{(X,\Delta)}.$ The corresponding constant $Z_{N,0}$ shall then be called the $\mathcal{K}$-reduced partition function, in view of its statistical mechanical interpretation described in Section 6.2. The fixed metric on $K_{(X,\Delta)}$ also determines the Mabuchi functional $\mathcal{M}$ on the space of Kähler metrics in $c_{1}(-K_{(X,\Delta)})$ (which, a priori, is only defined up to an additive constant).
@@
(1.16) $\text{(i)}-\lim_{N\rightarrow\infty}N^{-1}\log Z_{N,0}=\inf\mathcal{M}.$
As a consequence, $Z_{N,0}<\infty$ for $N$ sufficiently large and viewing the empirical measure $\delta_{N}$ as a random variable on $\left(X^{N},\mu_{0}^{(N)}\right),$
@@
(1.18) $(\delta_{N})_{*}\nu_{0}^{(N)}\sim e^{-NF(\mu)},$
where $F$ is the free energy functional, introduced in [11], that may be identified with the Mabuchi functional on the space of all Kähler metrics $\omega$ in $c_{1}(-K_{(X,\Delta)})$ via the relation $F(\omega^{n}/V)=\mathcal{M}(\omega)$. We show that the validity of such a LDP would imply both the asymptotics 1.16 and the convergence 1.17, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$. Moreover, by introducing temperature into the picture the corresponding LDP would also imply Conjecture 1.2, using the observation that the infimum over $\mathcal{G}$ in formula 1.8 is attained at the unique Kähler metric in a given $\mathcal{G}$–orbit whose volume form has vanishing moment (see Prop 3.4).
Our next main result establishes the first part of the previous conjecture on log Fano curves with an explicit rate of convergence:
@@
The convergence 1.16 holds on any K-polystable log Fano curve $(X,\Delta)$ when $N$ is taken as the sequence of odd positive integers. More precisely,
$N^{-1}\log Z_{N,0}=\inf\mathcal{M}+O\left(\frac{\log N}{N}\right).$
As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.17) we will show, in the companion paper [5], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 6.5).
The assumption that $N$ be odd in the previous theorem is equivalent to 0 being a regular value of $\boldsymbol{m}_{N}.$ In particular, this ensures that the level set $\left\{\boldsymbol{m}_{N}=0\right\}$ is a smooth submanifold of $X^{N}.$ To bypass this assumption one can, for a general log Fano manifold $(X,\Delta),$ replace $\left\{\boldsymbol{m}_{N}=0\right\}$ with a neighbourhood $\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon_{N}\right\}$ for an appropriate sequence $\epsilon_{N}$ tending to zero. The assumption that $(X,\Delta)$ be strongly uniformly Gibbs polystable in Theorem 1.4 implies that the corresponding normalizing constant $Z_{N,\epsilon_{N}}$ is finite, when $N\gg 1$ and $\epsilon_{N}\gg N^{-1/2}.$ Under this assumption we establish in Theorem 7.4 an effective large deviation lower bound on $Z_{N,\epsilon_{N}},$ which implies both Theorem 1.4 and the bound
$\limsup_{N\rightarrow\infty}-N^{-1}\log Z_{N,\epsilon_{N}}\leq\inf\mathcal{M},$
assuming that the Futaki character of $(X,\Delta)$ vanishes.
@@
where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{0}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.19 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [76]. Such integrals can be computed explicitly in terms of Gamma values, as first shown in [69] in the context of Conformal Field Theory, and letting $N\rightarrow\infty$ allows us to deduce Theorem 1.9, by exploiting some connections to K-stable arithmetic log Fano curves in [4].
It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.19 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 9, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [4], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [4], is a logarithmic generalization of Odaka’s modular height [104]. The proof of Theorem 1.9 reveals that this arithmetic conjecture holds for all K-polystable log Fano curves $(X,\Delta)$ defined over the integers.
@@ 1.7. Outlook on valuative aspects
where $X^{\text{div}}$ denotes the space of all divisorial valuations $v$ over $X,$ $A_{(X,\Delta)}(v)$ denotes the log discrepancy of $v$ (formula 2.8), $\left\|\mu\right\|$ denotes the non-Archimedean pluricomplex energy of a probability measure $\mu$ on $X^{\text{div}}$ with respect to $K_{(X,\Delta)}$ [41] and $\mathcal{P}(X^{\text{div}})_{\text{ss}}$ denotes the space of all probability measures $\mu$ on $X^{\text{div}}$ of the following form (where $N$ is the sequence defined by formula 1.6):
@@
The convergence in the second item above is motivated by the approach for proving the convergence of $\gamma^{(N)}(X,\Delta)$ towards $\delta(X,\Delta)_{\text{ }}$ outlined in [18, Section 5], using non-Archimedean pluripotential theory. The validity of the third item would imply the first item, by [61]. It should be emphasized that, in general, the invariant $\delta(X,\Delta)_{\text{ps }}$ differs from the reduced $\delta$-invariant of $(X,\Delta),$ introduced in [111], which is defined with respect to a fixed maximal complex torus in $\mathcal{G}.$ Our results imply Conjecture 1.10 for log Fano curves, but we leave the proof and further investigations of Conjecture 1.10 for the future.
@@ 1.8. Acknowledgements
Thanks to Sébastien Boucksom and Mattias Jonsson for discussions. This work was supported by a Wallenberg Scholar grant from the Knut and Alice Wallenberg foundation.
@@ 1.9. Organization of the paper
  • In Section 4 applications of Theorem 3.13 to constrained logarithmic Hardy–Littlewood–Sobolev (HLS) inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
  • In Section 5 we establish a sharp version of the constrained logarithmic HLS inequalities on the two-sphere $S^{2},$ by combining Theorem 3.13 with the uniqueness result for the Liouville equation on $S^{2}$ established in [80]. Applications to quantiative stability results for the logarithmic HLS inequality on $S^{2}$ are then given.
  • In Section 6 we study the $\mathcal{K}$–reduced probability measures $\mu_{0}^{(N)}$ on $X^{N}$ and the corresponding partition function $Z_{N,0}$ appearing in Conjecture 1.8. The asymptotics as $N\rightarrow\infty$ are described in a overarching conjectural Large Deviation Principle, formulated in Conjecture 6.7, that we show implies both Conjecture 1.1 and Conjecture 1.2.
  • In Section 7 we prove an effective large deviation bound for the thickened partition functions that implies Theorem 1.4
  • In Section 8 Theorem 1.9 is established, showing that $N^{-1}\log Z_{N,0}$ converges towards the inf of the Mabuchi functional on any log Fano curve with vanishing Futaki invariant.
@@ 2.1. Setup
We start by describing in more detail the general setup, introduced in Section 2. Let $(X,\Delta)$ be an $n$-dimensional log Fano manifold, i.e. $X$ is a complex projective manifold and $\Delta$ is an effective $\mathbb{Q}$–divisor on $X$ such that the log anticanonical line bundle $K_{(X,\Delta)}$ of $(X,\Delta)$ is ample. Recall that $K_{(X,\Delta)}$ is the $\mathbb{Q}$–line bundle defined as the tensor power of the canonical line bundle $K_{X}$ of $X$ and the $\mathbb{Q}$–line bundle defined by $\Delta,$ i.e. $K_{(X,\Delta)}\coloneqq K_{X}+\Delta,$ using the standard additive notation for tensor powers of line bundles that we shall adopt. Fix a positive integer $k$ such that $k\Delta$ is a divisor – rather than a mere $\mathbb{Q}$–divisor – so that $kK_{(X,\Delta)}$ defines a line bundle and set
@@
(2.1) $\text{Supp}\left(\mathcal{D}^{(N)}\right)\coloneqq\left\{(x_{1},\dots,x_{N})\colon\exists s\in H^{0}(X,-kK_{(X,\Delta)})-\{0\}\text{ such that }s(x_{1})=\cdots=s(x_{N})\right\}$